# OceanSim: A GPU-Accelerated Underwater Robot Perception Simulation Framework

### v1

@@ Abstract
Underwater simulators offer support for building robust underwater perception solutions. Significant work has recently been done to develop new simulators and to advance the performance of existing underwater simulators. Still, there remains room for improvement on physics-based underwater sensor modeling and rendering efficiency. In this paper, we propose OceanSim, a high-fidelity GPU-accelerated underwater simulator to address this research gap. We propose advanced physics-based rendering techniques to reduce the sim-to-real gap for underwater image simulation. We develop OceanSim to fully leverage the computing advantages of GPUs and achieve real-time imaging sonar rendering and fast synthetic data generation. We evaluate the capabilities and realism of OceanSim using real-world data to provide qualitative and quantitative results. Code and detailed documentation will be released to benefit the marine robotics community.
@@ II-A Underwater Simulators
In recent years, game engine-powered simulators (HoloOcean[6], UNav-Sim[8]) have achieved improved rendering quality by leveraging Unreal Engine [16]. Still, the official release version of HoloOcean [6] relies on Unreal Engine 4, which does not support rendering depth images. This limits the possibility of applying physics-based models for realistic underwater image rendering. HoloOcean also lacks GPU-based sonar rendering, which leads to slow rendering time for synthetic sonar data generation. Furthermore, HoloOcean requires building an Octree cache for sonar rendering, which takes extended time when a new environment is loaded. UNav-Sim [8] employs a newer version of Unreal Engine [16], UE5.1, to achieve more realistic underwater image rendering. However, the exact algorithm of the image formation model is inaccessible as Unreal Engine is closed-source. In addition, there is no example underwater scene configuration provided in UNav-Sim, making it more difficult for users to construct an underwater digital twin. Furthermore, UNav-Sim does not support simulating acoustic sensors such as imaging sonars, which are important for underwater perception. Therefore, there still remains a gap in development of a high-performance underwater robot perception simulator that can efficiently generate high-fidelity multi-modal data.
@@ II-B NVIDIA Isaac Sim
NVIDIA Isaac Sim [17] is a high-performance simulator designed for GPU-accelerated robotics simulation with superior physics accuracy and efficient tools for generating synthetic data. NVIDIA Isaac Sim has enabled successful digital twin construction and sim-to-real transfer in a wide range of fields [26, 27, 28, 29, 30, 19, 20, 31, 21, 22, 23]. Key features of NVIDIA Issac Sim include physics-based photorealistic rendering, seamless integration with NVIDIA Omniverse [18] and OpenUSD [32] ecosystems, developer-friendly 3D workflows, and GPU-based ray tracing. Recently, MarineGym has extended NVIDIA Issac Sim to the underwater domain [33]. MarineGym focuses on characterizing the simulation capability of NVIDIA Issac Sim for robot controls tasks and does not provide open-source implementation. On the other hand, our proposed simulation framework, OceanSim, focuses on building a high-performance simulator for underwater robot perception, and contributes novel implementations of common underwater sensors in NVIDIA Isaac Sim. OceanSim will be released as open-source to support the underwater robotics research community.
@@ III-A1 Camera
$I_{c}=Je^{-\beta_{\text{attn,c}}}+B_{\infty,c}\left(1-e^{-\beta_{\text{bs,c}}d}\right),$ (1)
@@ III-A2 Imaging Sonar
The rendering of OceanSim’s imaging sonar follows that of HoloOcean [7]. After obtaining the set of queried returns $\mathcal{Q}=\{x_{i}\in\mathbb{R}^{3}|\ x_{i}\text{ in view and not occluded}\}$ where $x_{i}$ is a point in the scene, we bin all the returns by range and azimuth. Then, we add additive and multiplicative noise to model the speckle noise properties of real imaging sonar. We begin with the additive noise term $\tilde{w}^{sa}$:
@@
We set up the default parameters of our proposed sonar model based on a Blueprint Subsea Oculus M750-d sonar, which is a common multibeam imaging sonar used in prior works [39, 40, 41, 42]. The Oculus M750-d used in our experiments has a 130 horizontal field-of-view and a 20 vertical field of view. To emulate its behavior to the best extent, we perform a range-wise normalization after collapsing the 3D intensity data to a 2D polar grid. We provide user-friendly APIs to choose between various plot styles, binning, and normalization methods.
@@ IV-A Demonstration of High-fidelity Digital Twin
Figure 6: Qualitative comparison of real-world data (top) and the simulated sonar-image pair (bottom).
@@ IV-B Quantitative Comparison of Sonar Rendering Efficiency
One of the core contributions of OceanSim is the accelerated rendering speed for sonar images, which is essential for generating large amounts of simulated sonar data to support training for data-driven perception models. To validate this feature, we quasonar ntitatively compare the rendering speed of OceanSim with Ho.
To ensure a fair comparison of time cost, all experiments on both HoloOcean and OceanSim are conducted on the same desktop, which has an NVIDIA RTX A6000 GPU, AMD Ryzen 5950X CPU, and $128$G RAM. We also unify the scenes and sensor configurations in both simulators. It is worth mentioning that we are not able to export the default simulation environments used in HoloOcean [6]. Therefore, we follow the documentation to import the scenes selected for comparison into HoloOcean using the Unreal Engine editor. We select three example scenes, including the MHL towing tank, a pier, and a simulated underwater mall scene. In practice, to quantify the rendering speed, we place the sensor rig at the same location in both simulators and tick the simulator to generate $100$ frames of sensor measurements. As discussed in Sec. II-A, HoloOcean [6] requires building a partial octree on first startup of a new scene. We also report the time needed to build the cache of the partial octree. In the follow-up sensor data generation, since the sensor rig is kept static, we ensure the quantified rendering speed of HoloOcean directly represents its best performance with pre-built cache.
Table II demonstrates that OceanSim significantly outperforms HoloOcean in sonar rendering speed. Notably, OceanSim benefits from GPU-based ray tracing to achieve real-time sonar rendering speed without requiring extra time for building cache. This feature is important to enabling large-scale simulated data generation for underwater robotics applications.

### v2

@@ Abstract
Underwater simulators offer support for building robust underwater perception solutions. Significant work has recently been done to develop new simulators and to advance the performance of existing underwater simulators. Still, there remains room for improvement on physics-based underwater sensor modeling and rendering efficiency. In this paper, we propose OceanSim, a high-fidelity GPU-accelerated underwater simulator to address this research gap. We propose advanced physics-based rendering techniques to reduce the sim-to-real gap for underwater image simulation. We develop OceanSim to fully leverage the computing advantages of GPUs and achieve real-time imaging sonar rendering and fast synthetic data generation. We evaluate the capabilities and realism of OceanSim using real-world data to provide qualitative and quantitative results. The code and detailed documentation are made available on the project website to support the marine robotics community: https://umfieldrobotics.github.io/OceanSim.
@@ II-A Underwater Simulators
In recent years, game engine-powered simulators (HoloOcean[6], UNav-Sim[8]) have achieved improved rendering quality by leveraging Unreal Engine [16]. Still, the official release version of HoloOcean [6] relies on Unreal Engine 4, which does not support rendering depth images. This limits the possibility of applying physics-based models for realistic underwater image rendering. HoloOcean also lacks GPU-based sonar rendering, which leads to slow rendering time for synthetic sonar data generation. Furthermore, HoloOcean requires building an Octree cache for sonar rendering, which takes extended time when a new environment is loaded. UNav-Sim [8] employs a newer version of Unreal Engine [16], UE5.1, to achieve more realistic underwater image rendering. However, the image formation model is difficult to customize and has not been validated against real underwater imagery. In addition, there is no example underwater scene configuration provided in UNav-Sim, making it more difficult for users to construct an underwater digital twin. Furthermore, UNav-Sim does not support simulating acoustic sensors such as imaging sonars, which are important for underwater perception. Therefore, there still remains a gap in development of a high-performance underwater robot perception simulator that can efficiently generate high-fidelity multi-modal data.
@@ II-B NVIDIA Isaac Sim
NVIDIA Isaac Sim [17] is a high-performance simulator designed for GPU-accelerated robotics simulation with superior physics accuracy and efficient tools for generating synthetic data. NVIDIA Isaac Sim has enabled successful digital twin construction and sim-to-real transfer in a wide range of fields [26, 27, 28, 29, 30, 19, 20, 31, 21, 22, 23, 32]. Key features of NVIDIA Issac Sim include physics-based photorealistic rendering, seamless integration with NVIDIA Omniverse [18] and OpenUSD [33] ecosystems, developer-friendly 3D workflows, and GPU-based ray tracing. Recently, MarineGym has extended NVIDIA Issac Sim to the underwater domain [34]. MarineGym focuses on characterizing the simulation capability of NVIDIA Issac Sim for robot controls tasks and does not provide open-source implementation. On the other hand, our proposed simulation framework, OceanSim, focuses on building a high-performance simulator for underwater robot perception, and contributes novel implementations of common underwater sensors in NVIDIA Isaac Sim. OceanSim will be released as open-source to support the underwater robotics research community.
@@ III-A1 Camera
$I_{c}=Je^{-\beta_{\text{attn,c}}d}+B_{\infty,c}\left(1-e^{-\beta_{\text{bs,c}}d}\right),$ (1)
@@ III-A2 Imaging Sonar
The rendering of OceanSim’s imaging sonar follows that of HoloOcean [7]. After obtaining the set of queried returns $\mathcal{Q}=\{x_{i}\in\mathbb{R}^{3}|\ x_{i}\text{ in view and not occluded}\}$ where $x_{i}$ is a point in the scene, we bin all the returns on a range-azimuth polar grid by summing up intensity measurements that fall into each grid. Then, we add additive and multiplicative noise to the binned value to model the speckle noise properties of real imaging sonar. We begin with the additive noise term $\tilde{w}^{sa}$:
@@
We set up the default parameters of our proposed sonar model based on a Blueprint Subsea Oculus M750-d sonar, which is a common multibeam imaging sonar used in prior works [40, 41, 42, 43]. The Oculus M750-d used in our experiments has a 130 horizontal field-of-view and a 20 vertical field of view. To emulate its behavior to the best extent, we perform a range-wise normalization after collapsing the 3D intensity data onto the 2D grid. We provide user-friendly APIs to choose between various plot styles, binning, and normalization methods.
@@ IV-A Demonstration of High-fidelity Digital Twin
Figure 6: Qualitative comparison of real-world data (top) and the simulated sonar-image pair (bottom). For the above simulated sonar image, we use 3000 x 460 for the raycast density at horizontal and vertical FOVs, and 350 x 220 for the number of bins in range and azimuth directions.
@@ IV-B Quantitative Comparison of Sonar Rendering Efficiency
One of the core contributions of OceanSim is the accelerated rendering speed for sonar images, which is essential for generating large amounts of simulated sonar data to support training for data-driven perception models. To validate this feature, we quantitatively compare the sonar rendering speed of OceanSim with HoloOcean [6].
To ensure a fair comparison of time cost, all experiments on both HoloOcean and OceanSim are conducted on the same desktop, which has an NVIDIA RTX A6000 GPU, AMD Ryzen 5950X CPU, and 128128G RAM. We also unify the scenes and sensor configurations in both simulators. It is worth mentioning that we are not able to export the default simulation environments used in HoloOcean [6]. Therefore, we follow the documentation to import the scenes selected for comparison into HoloOcean using the Unreal Engine editor. We select three example scenes, including the MHL towing tank, a pier, and a simulated underwater mall scene. In practice, to quantify the rendering speed, we place the sensor rig at the same location in both simulators and tick the simulator to generate $100$ frames of sensor measurements. As discussed in Sec. II-A, HoloOcean [6] requires building a partial octree on first startup of a new scene. We also report the time needed to build the cache of the partial octree. In the follow-up sensor data generation, since the sensor rig is kept static, we ensure the quantified rendering speed of HoloOcean directly represents its best performance with pre-built cache.
Table II demonstrates that OceanSim significantly outperforms HoloOcean in sonar rendering speed. Notably, OceanSim benefits from GPU-based ray tracing to achieve real-time sonar rendering speed without requiring extra time for building cache. This feature is important to enabling large-scale simulated data generation for underwater robotics applications. We also observe that the rendering speed is more impacted by scaling up the resolution (i.e. the number of bins) versus increasing the raycast density, and overly increasing resolutions can lead to fidelity drop due to the rendered images appearing much sharper than the real world ones.